Determine own price elasticity of demand if inverse demand is 1)P=1/(3+X)^2
X=2/8+P
X=P^4/3
1
Expert's answer
2019-06-12T10:55:01-0400

If P=1/(3+X)^2, then

"X = |1\/P^{0.5}| - 3, Ed = -0.5\u00d7P\/Q."

If X=2/8+P, then Ed = P/Q.

If X=P^4/3, then Ed = 4/3P/Q.